OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 3x + y - z = 2 2x + y + 2z = 8 2x - 3y + z = 21 A. (3, 4, 3) B. (-3, 4, 3) C. (3, -4, 3) D. (3, 4, -3)

OpenStudy (anonymous):

@texaschic101

OpenStudy (john_es):

C is the solution. Do you understand why?

OpenStudy (anonymous):

3x + y - z = 2 2x - 3y + z = 21 5x - 2y = 23 3x + y - z = 2 2x + y + 2z = 8 3x - 2x + y - y - z - 2z = 2 - 8 x - 3z = -6 2x - 2x + y + 3y + 2z - z = 8 - 21 4y + z = -13 z = -13 - 4y x = -6 + 3z = -6 + 3 * (-13 - 4y) = -6 - 39 - 12y = -45 - 12y 3x + y - z = 2 3 * (-45 - 12y) + y - (-13 - 4y) = 2 -135 - 36y + y + 13 + 4y = 2 5y - 36y - 122 = 2 -31y = 124 y = -4 z = -13 - 4y z = -13 + 16 z = 3 x = -6 + 3z x = -6 + 9 x = 3 (3 , -4 , 3)

OpenStudy (anonymous):

so yes its C